OpenStudy (anonymous):

Use the Law of Levers to solve the problem. Three friends pay $100 each to share a cottage on a camping trip. The cost per person varies inversely as the number of people on the camping trip. How many people would have to be on the trip to make the cost $60 per person?

OpenStudy (anonymous):

p = price per person k = some constant n = number of people first solve for k 100=(k/3) k=300 then solve for new n 60=(300)/n n=5